The average weight of a professional football player in 2009 was 255.6 pounds. Assume the population standard deviation is 30The answers are in red and thanks in advance
Answer #1

POX-29- La 3013) Q. @ 52=0/16 = 307531 - 15.39) Opaž<242) Plze 242-255.6) - PIZZ-2.52) 210.0059) © Pl#>240)- P(z>246-255.6 )Some answers are a bit different in decimal because you might have used 5.39 for all other problems, but I’ve taken 30/310.5 everywhere.